# Elastic Full-Waveform Inversion : How the physics of problem improves data-driven techniques?

### v1

@@ Abstract
Full-Waveform Inversion (FWI) is a nonlinear iterative seismic imaging technique that, by reducing the misfit between recorded and predicted seismic waveforms, can produce detailed estimates of subsurface geophysical properties. Nevertheless, the strong nonlinearity of FWI can trap the optimization in local minima. This issue arises due to factors such as improper initial values, the absence of low frequencies in the measurements, noise, and other related considerations. To address this challenge and with the advent of advanced machine-learning techniques, data-driven methods, such as deep learning, have attracted significantly increasing attention in the geophysical community. Furthermore, the elastic wave equation should be included in FWI to represent elastic effects accurately. The intersection of data-driven techniques and elastic scattering theories presents opportunities and challenges. In this paper, by using the knowledge of elastic scattering (Physics of problem) and integrating it with deep learning techniques, we propose methods for the solution of time-harmonic FWI to enhance accuracy compared to pure data-driven approaches. Moreover, by modifying the structure of the Variational Autoencoder, we introduce a probabilistic deep learning method based on the physics of the problem that enables us to explore the uncertainties of the solution. According to the limited availability of datasets in this field and to assess the performance and accuracy of the proposed methods, we create a comprehensive dataset close to reality and conduct a comparative analysis of the presented approaches to it.
@@
Full Waveform Inversion - Inverse Elastic Scattering - Data-Driven Technique - Deep Learning - Variational Autoencoder - Seismic Imaging
@@ 1 Introduction
Data-driven methods, particularly deep learning techniques, have shown a remarkable capacity to approximate nonlinear mapping functions between different data domains, including natural language processing [40], speech recognition [41], and many other fields [42, 43], especially for inverse problems like image reconstruction [44], image super-resolution [45, 46], etc. These cutting-edge innovations offer new possibilities for Inverse Scattering Problem (ISP) and consequently seismic imaging. Deep learning approaches provide a forthcoming imaging technique by avoiding bottlenecks mentioned earlier [47, 48]. One main topic in this research area is to build a direct inverse mapping from observations to subsurface structure by training neural networks on paired data of seismic waveforms and desired parameters of the model [49, 50, 51, 52]. Actually, instead of solving the wave equation, this method considers FWI as a simple deep-learning problem, the same as image recognition. Some research efforts have been conducted in this field, such as VAE, Unet, and RNN, as detailed in the comprehensive review article [53]. The techniques as mentioned above, can achieve exceptional efficiency once fully trained. However, the generalizability of data-driven approaches is constrained by the size and diversity of the training set. Furthermore, the availability of data is relatively limited in this field due to the high acquisition cost. According to the high commercial value of this data, a very limited amount of it is accessible to the public. Moreover, since these models learn from data directly without incorporating domain knowledge, they are susceptible to generating unreasonable or unrealistic predictions that may not adapt to the physical mechanism. As a result, these observations have prompted researchers to integrate physics-based and data-driven methods.
@@
The simultaneous use of the physics-based method and data-driven approach is an idea that has recently gained attention, which is called the Physics-Guided Machine Learning (PGML) approach. Numerous studies have shown how crucial prior knowledge is to enhancing the effectiveness of the physics-based FWI optimization technique. The existing parameter distribution [54, 55, 56], the low-frequency component of the measured signals [57, 58], assistance in resolving the cycle skipping problem, and other things are examples of this prior knowledge. On the other hand, Physics-Informed Neural Networks (PINNs) [59, 60] have become an effective tool for resolving inverse problems, mainly when insufficient scatter measurements are available, and no complete system knowledge is known. PINNs attempt to incorporate physics into the trained network, commonly achieved by augmenting the loss function with a residual term derived from the governing equation. This approach effectively regularizes the solution space during training. Specifically, incorporating the wave equation in data-driven methods has been associated with some factors, such as improved network accuracy, enhanced generalization, increased robustness, a reduction in the amount of required training data, and, in certain instances, the elimination of the need for labeled data. Like PINN, several works have been completed by applying deep learning and PDEs simultaneously to seismic FWI [61, 62, 63]. Estimating the parameters of interest using a learned representation of the underlying distribution of the physical parameters is an impressive method of integrating physics into deep learning. More specifically, a widely used network architecture utilized to describe data distribution is the generative adversarial network (GAN). This category is illustrated by several recent studies [64, 65]. Other methods within PGML include Data Augmentation [66, 67], Unrolling-Based Techniques, which help to reduce computation efficiently [68], and the Learning Loss Function, which chooses an appropriate loss function [69, 70]. A comprehensive overview of works in this field is provided in the article [71].
@@
This paper introduces approaches to solving time-harmonic elastic FWI through data-driven techniques and their combination within physical constraints. We subsequently conduct a comparative analysis of these methods. A consequence of the inherent ill-conditioning of the problem is that small noises in the data can lead to significant errors, which is evident in the proposed methods. Some tricks are employed to mitigate this issue. Another challenge arises from the fact that numerous plausible reconstructions may fit with the noisy measurements. This variety of solutions recommends employing techniques that recover multiple reconstructions. A probabilistic framework in characterizing solutions enhances the reliability of interpretation for reconstructions, particularly when the problem becomes more ill-posed. In this paper, we introduce a novel probabilistic method based on Variational Autoencoder (VAE). In all the methods to be discussed, we make the assumption that the Lame parameters are constant and known. Additionally, we assume that the receivers have collected information (displacement field) on the surface of the field. From this collected data, we want to acquire the density of the field in the entire desired area.
@@
Section 2 contains the physics of the problem and the relations between variables. To use data-driven techniques to solve the problem, we create a dataset. In section 6, we aim to produce more complex and comprehensive models by making artificial fields. These artificial fields can be good simulators for natural fields. So, we use them and their related displacement field (derived from the governed elastic wave equation) in the learning process. In the first method 3, which is a pure data-driven approach, we employ a Convolutional Neural Network (CNN) to solve the inverse problem. This CNN takes surface information as input and produces the density of the subsurface as output. In the second method 4, which is a PGML approach, we utilize CNN to derive the displacement field over the entire area from the information on the surface. Subsequently, using different techniques and leveraging the physics of the problem, we determine the desired density. The third method 5, which is a probabilistic deep learning inverse approach, inspired by the structure of VAE, we aim to introduce a method that not only generates solutions but also can provide a probabilistic solution for the inverse problem. Additionally, this method can utilize the physics of the problem to enhance and refine its results.
@@ 2 Governing equations and main results
$\mathbf{G}\big{(}\boldsymbol{x},\boldsymbol{y},\omega\big{)}=\dfrac{1}{\mu}\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{s}\big{)}\mathbf{I}+\dfrac{1}{\omega^{2}}\nabla_{x}\nabla_{x}^{\perp}\Big{(}\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{s}\big{)}-\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{p}\big{)}\Big{)}$
@@
The details and proof of this theorem are given in [7, 8, 9]. In the following sections, we will explain the methodology for elastic FWI, which includes data-driven techniques and the integration of principles of physics related to the problem with it. Our goal is to recover the subsurface using the data we have collected from the surface of the field. To achieve our objective, we have created a synthetic dataset, which is thoroughly described in section 6 using equation 4. We divided the region $\Omega$ into $N=n\times n$ parts, assuming that we have the displacement field $\boldsymbol{u}=(u_{1},u_{2})$ at the Earth’s surface, with $2n$ values for each incident wave. We aim to determine the density $\rho$, which consists of $N$ unknown values.
@@ 4 Second method
In this section, initially, we train a CNN to obtain the displacement field $\boldsymbol{u}$ within region $\Omega$. This network takes the same input as the one used in method 3 and provides the displacement fields within region $\Omega$ as its output. These displacement fields are collectively denoted as $\overline{U}=\{\overline{U}_{1},...,\overline{U}_{k}\}$ and correspond to all $k$ incident waves.
@@
Using the obtained displacement field and the integral equation 4, we want to find the density of the desired area. We aim to simultaneously apply deep learning and the problem’s physics to improve accuracy through some techniques. It should be noted that in this section, for convenience, we use $\rho$ instead of $1-\rho$ in formula 4.
@@ 4.1 Least Square
Considering that we have the displacement field across the entire desired area and the use of relation 4, the initial approach involves solving the problem using the least squares method. So, it is enough to collect all the information and write it as a linear system.
@@
A suitable technique for solving the equation mentioned above is the least-squares method with $L_{2}$  regularization. One notable challenge when dealing with equation 8 is its considerable size. With an increment in the number of incident waves and $N$, which leads to an increase in matrix size, using this method can become quite challenging. The main difficulty in this method is that matrix $A$ is ill-posed in relation 6. When this matrix is combined with the displacement vector $\overline{U}$, which may contain a small error, this ill-posedness leads to a significant error in the solution of the problem.
@@ 4.2 Inverse Convolution
The integral form of relation 4 and the Green’s tensor $\mathbf{G}$ in it, evokes a convolution-like relation in our mind. By defining this particular type of convolution, we want to obtain the inverse of $\mathbf{G}$ with this convolution.
@@
Note that in relation 12, we drive four approximated densities. Furthermore, for each $\boldsymbol{u}^{inc}$ and its corresponding $\overline{\boldsymbol{u}}$, we can form the relation 12 to obtain $\rho$. While each of these densities can be utilized to estimate the original density, deciding which one to choose is challenging. It is crucial to have a framework that can provide a unique estimate by incorporating these approximations. Employing the Unet network structure presents a promising solution. Since we will encounter a similar situation in the upcoming subsection, we will elaborate on the application of the Unet structure and provide more comprehensive details on this problem in method 4.3.
@@ 4.3 Linear-to-Nonlinear
where $U_{j},U^{inc}_{j}$ and $,\hat{\rho}$ are the same as defined in relation 6. Our objective is to find a structure for the $\hat{G}$ function that not only guarantees high accuracy but also avoids a substantial computational cost. Furthermore, this function should also has a robust performance in the presence of noisy data and mitigate their impact. We define $\hat{G}$ function as a nonlinear combination of $\hat{G_{i}}\,,\,\,i=1,...,m$ convolutions:
@@
A clever technique to obtain functions $\hat{G_{i}}\,,\,\,i=1,...,m$ is to train them using the available dataset. If we consider the relationship described above as a network, its input would be $U-U^{inc}$, and its output would be $\hat{\rho}U$. To enable the network to effectively denoise, we add Gaussian noises to the input of this network, which are $U-U^{inc}$ functions, and we expect it to give us the exact $\hat{\rho}U$ function as output. It is important to note that during the training of this network, the original $U$ is used as an input. However, when we apply the network in practice, we replace it with an estimated $\overline{U}$. We expect the difference between $U$ and $\overline{U}$ to be a Gaussian random field. By adding a Gaussian random field to the input, we aim to avoid noise and ensure the network understands the nature of the true field. This process reduces the errors resulting from the approximation of $U$ in the system and enhances its stability. The structure of this network is illustrated in Figure 3.
@@
So far, we have followed a two-step process. Initially, in section 4, we obtained an estimate for $U$. Subsequently, we employed the network constructed based on relations 13 and 14 to estimate the values of $\hat{\rho}U$ (which we have called $\overline{\rho U}$). Consequently, the density is computed by dividing $\overline{\rho U}$ by $\overline{U}$. However, it is worth noting that the number of these approximated densities amounts to $4k$, as we have $k$ incident waves, and for each of them, there exist horizontal, vertical, real, and imaginary components.

### v2

# Integrating Physics of the Problem into Data-Driven Methods to Enhance Elastic Full-Waveform Inversion with Uncertainty Quantification

@@ Abstract
Full-Waveform Inversion (FWI) is a nonlinear iterative seismic imaging technique that, by reducing the misfit between recorded and predicted seismic waveforms, can produce detailed estimates of subsurface geophysical properties. Nevertheless, the strong nonlinearity of FWI can trap the optimization in local minima. This issue arises due to factors such as improper initial values, the absence of low frequencies in the measurements, noise, and other related considerations. To address this challenge and with the advent of advanced machine-learning techniques, data-driven methods, such as deep learning, have attracted significantly increasing attention in the geophysical community. Furthermore, the elastic wave equation should be included in FWI to represent elastic effects accurately. The intersection of data-driven techniques and elastic scattering theories presents opportunities and challenges. In this paper, by using the knowledge of elastic scattering (physics of the problem) and integrating it with machine learning techniques, we propose methods for the solution of time-harmonic FWI to enhance accuracy compared to pure data-driven and physics-based approaches. Moreover, to address uncertainty quantification, by modifying the structure of the Variational Autoencoder, we introduce a probabilistic deep learning method based on the physics of the problem that enables us to explore the uncertainties of the solution. According to the limited availability of datasets in this field and to assess the performance and accuracy of the proposed methods, we create a comprehensive dataset close to reality and conduct a comparative analysis of the presented approaches to it.
@@
Full Waveform Inversion - Inverse Elastic Scattering - Data-Driven Technique - Deep Learning - Variational Autoencoder - Seismic Imaging - Uncertainty Quantification
@@ 1 Introduction
Data-driven methods, particularly deep learning techniques, have shown a remarkable capacity to approximate nonlinear mapping functions between different data domains, including natural language processing [40], speech recognition [41], and many other fields [42, 43], especially for inverse problems like image reconstruction [44], image super-resolution [45, 46], etc. These cutting-edge innovations offer new possibilities for Inverse Scattering Problem (ISP) and consequently seismic imaging. Deep learning approaches provide a forthcoming imaging technique by avoiding bottlenecks mentioned earlier [47, 48]. One main topic in this research area is to build a direct inverse mapping from observations to subsurface structure by training neural networks on paired data of seismic waveforms and desired parameters of the model [49, 50, 51, 52]. Actually, instead of solving the wave equation, this method considers FWI as a simple deep-learning problem, the same as image recognition. Some research efforts have been conducted in this field, such as VAE, GAN, Unet, and RNN, as detailed in the comprehensive review articles [53, 54]. The techniques as mentioned above, can achieve exceptional efficiency once fully trained. However, the generalizability of data-driven approaches is constrained by the size and diversity of the training set. Furthermore, the availability of data is relatively limited in this field due to the high acquisition cost. According to the high commercial value of this data, a very limited amount of it is accessible to the public. Moreover, since these models learn from data directly without incorporating domain knowledge, they are susceptible to generating unreasonable or unrealistic predictions that may not adapt to the physical mechanism. As a result, these observations have prompted researchers to integrate physics-based and data-driven methods.
@@
The simultaneous use of the physics-based method and data-driven approach is an idea that has recently gained attention, which is called the Physics-Guided Machine Learning (PGML) approach. Numerous studies have shown how crucial prior knowledge is to enhancing the effectiveness of the physics-based FWI optimization technique. The existing parameter distribution [55, 56, 57], the low-frequency component of the measured signals [58, 59], assistance in resolving the cycle skipping problem, and other things are examples of this prior knowledge. On the other hand, Physics-Informed Neural Networks (PINNs) [60, 61], a subset of PGML, become an effective tool for resolving inverse problems, mainly when insufficient scatter measurements are available, and no complete system knowledge is known. PINNs attempt to incorporate physics into the trained network, commonly achieved by augmenting the loss function with a residual term derived from the governing equation. This approach effectively regularizes the solution space during training. Specifically, incorporating the wave equation in data-driven methods has been associated with some factors, such as improved network accuracy, enhanced generalization, increased robustness, a reduction in the amount of required training data, and, in certain instances, the elimination of the need for labeled data [62, 63, 64]. Like PINN, several works have been completed by applying deep learning and PDEs simultaneously to seismic FWI [65, 66, 67, 68, 69]. Estimating the parameters of interest using a learned representation of the underlying distribution of the physical parameters is an impressive method of integrating physics into deep learning. More specifically, a widely used network architecture utilized to describe data distribution is the generative adversarial network (GAN). This category is illustrated by several recent studies [70, 71]. Other methods within PGML include Data Augmentation [72, 73], Unrolling-Based Techniques, which help to reduce computation efficiently [74], and the Learning Loss Function, which chooses an appropriate loss function [75, 76]. A comprehensive overview of works in this field is provided in the article [77].
Probabilistic methods within seismic inversion, including FWI and elastic FWI, allow for a more thorough exploration of the model space and quantifying uncertainties associated with the inversion results. These methods are typically implemented through Bayesian frameworks, where the posterior distribution of subsurface parameters is estimated, taking into account prior knowledge and observational data. Such methods can reveal the most likely model and provide a quantitative understanding of uncertainty within the inversion results. Among the relevant papers in this field, we can mention [78, 79, 80, 81]. In practice, these probabilistic methods often utilize Monte Carlo sampling, such as Hamiltonian Monte Carlo [82] and Markov Chain Monte Carlo (MCMC) [83, 84, 85] techniques, to explore the high-dimensional model space. Though these techniques approximate the posterior distribution, they are computationally intensive due to the requirement for numerous forward simulations. Some recent advancements aim to mitigate these computational demands by using techniques like low-rank approximations of the Hessian matrix and quasi-Newton methods, which reduce the number of necessary samples and computational resources [86, 87]. Advanced probabilistic methods, including ensemble-based approaches [88] and variational inference [89, 90], have also been explored, even though challenges related to computational cost and convergence remain significant hurdles. Additionally, machine learning and deep learning tools are increasingly being integrated with probabilistic FWI and elastic FWI, leveraging data-driven insights and reducing computational costs [91, 92, 55]. This evolving landscape not only enhances the accuracy of inversion results but also broadens the application of these methods in complex geological settings, making them valuable tools in the fields of seismic exploration and reservoir characterization. Furthermore, PINNs have attracted significant attention in probabilistic seismic inversion due to their ability to incorporate physics directly into the neural network training process. PINNs’ probabilistic formulation aids in capturing the inherent uncertainty in seismic inversion, yielding solutions that are both physically consistent and statistically reliable [93, 94].
@@
This paper introduces approaches to solving time-harmonic elastic FWI through data-driven techniques and their combination within physical constraints. We subsequently conduct a comparative analysis of these methods with pure data-driven and physics-based methods. A consequence of the inherent ill-conditioning of the problem is that small noises in the data can lead to significant errors, which is evident in the proposed methods. Some tricks are employed to mitigate this issue. Another challenge arises from the fact that numerous plausible reconstructions may fit with the noisy measurements. This variety of solutions recommends employing techniques that recover multiple reconstructions. A probabilistic framework in characterizing solutions enhances the reliability of interpretation for reconstructions, particularly when the problem becomes more ill-posed. In this paper, we introduce a novel probabilistic method based on Variational Autoencoder (VAE) for uncertainty quantification. In all the methods to be discussed, we make the assumption that the Lame parameters are constant and known. Additionally, we assume that the receivers have collected information (displacement field) on the surface of the field. From this collected data, we want to acquire the density of the field in the entire desired area.
@@
Section 2 contains the physics of the problem and the relations between variables. To use data-driven techniques to solve the problem, we create a dataset. In section 7, we aim to produce more complex and comprehensive models by making artificial fields. These artificial fields can be good simulators for natural fields. So, we use them and their related displacement field (derived from the governed elastic wave equation) in the learning process. In the first method 3, which is a pure data-driven approach, we employ a Convolutional Neural Network (CNN) to solve the inverse problem. This CNN takes surface information as input and produces the density of the subsurface as output. In the second method 4, which is a PGML approach, we utilize CNN to derive the displacement field over the entire area from the information on the surface. Subsequently, using different techniques and leveraging the physics of the problem, we determine the desired density. The third method 5, which is a probabilistic deep learning inverse approach, inspired by the structure of VAE, we aim to introduce a method that not only generates solutions but also can provide a probabilistic solution for uncertainty quantification. Additionally, this method can utilize the physics of the problem to enhance and refine its results. In the fourth method 6, two traditional physics-based techniques are implemented to facilitate comparison with the approaches presented in this paper.
@@ 2 Governing equations and main results
$\mathbf{G}\big{(}\boldsymbol{x},\boldsymbol{y},\omega\big{)}=\dfrac{1}{\mu}\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{s}\big{)}\mathbf{I}+\dfrac{1}{\omega^{2}}\nabla_{x}\nabla_{x}^{\perp}\Big{(}\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{s}\big{)}-\Phi\big{(}\boldsymbol{x},\boldsymbol{y},\kappa_{p}\big{)}\Big{)}$ (4)
@@
The details and proof of this theorem are given in [7, 8, 9].
Remark 2.2.
In the full waveform inversion problem, we usually seek three parameters: density, P-wave, and S-wave. The Green’s function introduced in equation 4, which is used in Theorem 2.1, is valid for constant Lame parameters. If we find a Green’s function for an elastic scattering problem with non-constant Lame parameters, either analytically or numerically, the results presented in this paper will be generalized. paper [96] suggests approaches to drive Green’s function for the heterogeneous elastic equations in specific cases, which can further generalize this study. Consequently, in this paper, we focus on finding the density and omit P-wave and S-wave.
In the following sections, we will explain the methodology for time-harmonic elastic FWI, which includes data-driven techniques and the integration of principles of physics related to the problem with it. Our goal is to recover the subsurface using the data we have collected from the surface of the field. To achieve our objective, we have created a synthetic dataset, which is thoroughly described in section 7 using equation 5. We divided the region $\Omega$ into $N=n\times n$ parts, assuming that the displacement field $\boldsymbol{u}=(u_{1},u_{2})$ is measured at $n$ receivers on the Earth’s surface, yielding to $2n$ values for each incident wave. Additionally, the number of incident waves (sources) is assumed to be $k$ with a single frequency for all of them.
@@ 4 Second method
As we know, the time-harmonic elastic FWI problem is nonlinear (in integral relation 5, $\rho$ and $\boldsymbol{u}$ are multiplied). If we have the values of $\boldsymbol{u}$ over the entire region $\Omega$, the problem becomes linear, although it is highly ill-conditioned. In this section, we will present methods for solving this problem by estimating $\boldsymbol{u}$ over the entire region.
In this section, initially, we train a CNN to obtain the displacement field $\boldsymbol{u}$ within region $\Omega$. This network takes the same input as the one used in the First method 3 and provides the displacement fields within region $\Omega$ as its output. These displacement fields are collectively denoted as $\overline{U}=\{\overline{U}_{1},...,\overline{U}_{k}\}$ and correspond to all $k$ incident waves.
@@
Using the obtained displacement field and the integral equation 5, we want to find the density of the desired area. We aim to simultaneously apply deep learning and the problem’s physics to improve accuracy through some techniques. It should be noted that in this section, for convenience, we use $\rho$ instead of $1-\rho$ in formula 5.
@@ 4.1 Least Square
Considering that we have the displacement field across the entire desired area and the use of relation 5, the initial approach involves solving the problem using the least squares method. So, it is enough to collect all the information and write it as a linear system.
@@
A suitable technique for solving the equation mentioned above is the least-squares method with $L_{2}$  regularization. One notable challenge when dealing with equation 9 is its considerable size. With an increment in the number of incident waves and $N$, which leads to an increase in matrix size, using this method can become quite challenging. The main difficulty in this method is that matrix $A$ is ill-posed in relation 7. When this matrix is combined with the displacement vector $\overline{U}$, which may contain a small error, this ill-posedness leads to a significant error in the solution of the problem.
@@ 4.2 Inverse Convolution
The integral form of relation 5 and the Green’s tensor $\mathbf{G}$ in it, evokes a convolution-like relation in our mind. By defining this particular type of convolution, we want to obtain the inverse of $\mathbf{G}$ with this convolution.
@@
Note that in relation 13, we drive four approximated densities. Furthermore, for each $\boldsymbol{u}^{inc}$ and its corresponding $\overline{\boldsymbol{u}}$, we can form the relation 13 to obtain $\rho$. While each of these densities can be utilized to estimate the original density, deciding which one to choose is challenging. It is crucial to have a framework that can provide a unique estimate by incorporating these approximations. Employing the Unet network structure presents a promising solution. Since we will encounter a similar situation in the upcoming subsection, we will elaborate on the application of the Unet structure and provide more comprehensive details on this problem in method 4.3.
@@ 4.3 Linear-to-Nonlinear
where $U_{j},U^{inc}_{j}$ and $\hat{\rho}$ are the same as defined in relation 7. Our objective is to find a structure for the $\hat{G}$ function that not only guarantees high accuracy but also avoids a substantial computational cost. Furthermore, this function should also has a robust performance in the presence of noisy data and mitigate their impact. We define $\hat{G}$ function as a nonlinear combination of $\hat{G_{i}}\,,\,\,i=1,...,m$ convolutions:
@@
A clever technique to obtain functions $\hat{G_{i}}\,,\,\,i=1,...,m$ is to train them using the available dataset. If we consider the relationship described above as a network, its input would be $U-U^{inc}$, and its output would be $\hat{\rho}U$. To enable the network to effectively denoise, we add Gaussian noises to the input of this network, which are $U-U^{inc}$ functions, and we expect it to give us the exact $\hat{\rho}U$ function as output. It is important to note that during the training of this network, the original $U$ is used as an input. However, when we apply the network in practice, we replace it with an estimated $\overline{U}$. We expect the difference between $U$ and $\overline{U}$ to be a Gaussian random field. By adding a Gaussian random field to the input, we aim to avoid noise and ensure the network understands the nature of the true wavefield $\hat{\rho}U$. This process reduces the errors resulting from the approximation of $U$ in the system and enhances its stability. The structure of this network is illustrated in Figure 3.
So far, we have followed a two-step process. Initially, in section 4, we obtained an estimate for $U$. Subsequently, we employed the network constructed based on relations 14 and 15 to estimate the values of $\hat{\rho}U$ (which we have called $\overline{\rho U}$). Consequently, the density is computed by dividing $\overline{\rho U}$ by $\overline{U}$. However, it is worth noting that the number of these approximated densities amounts to $4k$, as we have $k$ incident waves, and for each of them, there exist horizontal, vertical, real, and imaginary components.